# First Resolution of Microlensed Images of a Binary-Lens Event

### v1

@@ 4.3 Joint Analysis with Light Curves
In this section, we present joint analysis of VLTI and light-curve modeling. We begin by comparing the light-curve parameters and VLTI constraints. Figure 6 shows the comparisons between the $\alpha$ parameters $\alpha_{\rm LC}$ derived from the light-curve solutions with $\alpha_{\rm VLTI}$. The three solutions are mutually compatible within $3\,\sigma$, and they are all at $\Psi\sim-100^{\circ}$. These include one close solution (“close A--”), and two wide solutions (“wide A--” and “wide B$+$”). Next we check the consistency of the direction ${\bm{\pi}}_{\rm E}$ between the light curve solutions and VLTI solution at $\Psi\sim-100^{\circ}$. Figure 7 displays the posterior distribution of $(\pi_{\rm E,E},\pi_{\rm E,N})$ and the $\Phi_{{\bm{\pi}}}$ constraint from VLTI. For three light-curve solutions (“wide A--”, “wide B$+$” and “close A--”), the two directions agree to $\sim 1\,\sigma$.
@@
Then we perform a joint fit to the VLTI and light curve data for the “wide A--”, “wide B$+$” and “close A--” solutions. We include the higher-order effects of microlens parallax and binary orbital motion. Table 2 lists the best-fit parameters and their uncertainties. The global best-fit is the “wide A--” solution, and Figure 8 shows its best-fit closure-phase models along with the geometric configurations of the source, caustic and images. The “close A--” solution is worse by $\Delta\chi^{2}\approx 50$, which is somewhat ($\approx 15$) less compared with light-curve-only $\Delta\chi^{2}$. Therefore, the preference for the “wide A--” over the “close A--” comes from the light-curve fitting. It mainly stems from a $\sim 0.1$ mag broad bump-like feature in Gaia $G$-band light curve around $\rm HJD^{\prime}\sim 9300$, which can be only well matched by the wide solutions (see the right panel of Figure 4). In contrast, the “wide B$+$” model fit is worse by $\Delta\chi^{2}\approx 267$ with respect to “wide A--”, significantly increased by $\approx 218$ compared with light-curve only. As discussed in § 4.2, such $u_{0}>0$ geometries are strongly disfavored by the constraints from multi-epoch VLTI data.
@@
To further understand the VLTI constraints on these three solutions, we performed additional fitting with the interferometric data only. We hold the parameters $(s,q,\rho,\Psi,\theta_{\rm E})$, which define the caustics’ configuration, fixed at their best-fit values from the joint modeling, while the source positions $(u_{x},u_{y})$ of the three nights are allowed to vary. Figure 9 shows the $3\,\sigma$ $(u_{x},u_{y})$ contours (magenta) compared with the expected source positions (black dots) on the best-fit trajectory (black solid line) from the joint modeling for each solution. For the “wide A--” solution, there is a $\gtrsim 3\sigma$ discrepancy in $(u_{x},u_{y})$ during Night 1, which corresponds to the $\lesssim 0.3^{\circ}$ systematic residuals of the closure phases for U4U2U1 and U4U3U2 in both exposures of that night, as displayed in the top two sub-panels of Figure 8. During the other two nights, the source positions are in good agreements with the “wide A--” solution. Note that such $\sim 0.3^{\circ}$ systematic offsets also show up in the third exposure of Night 2 (see Figure 8), but they are not present in the other three exposures (i.e., Night 2-1, 2-2 and 2-4), which are in good agreement with the model. This suggests the existence of sub-degree level systematic error in the closure-phase data. The “close A--” solution shows a similar degree of consistency to the “wide A--”. In contrast, for the “wide B+” solution, which is ruled out by the VLTI data, the source positions of all three nights deviate by $>3\,\sigma$ from the expectation.
@@
In conclusion, because the “wide A--” solution is preferred over all other local minima by $\Delta\chi^{2}\gtrsim 50$, we regard it as the only favored solution of ASASSN-22av.
@@ 5.1 Source Properties and Finite-Source Effects
The finite-source effects exhibited in ASASSN-22av’s light curves provide an opportunity to cross check the VLTI-measured $\theta_{\rm E}$ with $\theta_{\rm E}={\theta_{*}}/{\rho}$. The standard method (Yoo et al., 2004) to derive $\theta_{*}$ for bulge microlensing events uses the source’s de-reddened color and magnitude with the extinction coefficients estimated from the location of bulge red clump (RC) stars in a color-magnitude diagram (CMD). In our case, the source star is not toward the bulge, and the CMD contains RC stars at a variety of distances, making it difficult to apply the standard method. Gaia DR3 reports a trigonometric parallax $\pi_{\rm S}=0.137\pm 0.016\,{\rm mas}$. We adopt $\pi_{{\rm S},\rm corr}=0.185\,{\rm mas}$ after applying the preliminary zero-point correction from Lindegren et al. (2021). The independent investigation on the Gaia parallax zero-point offset by Groenewegen (2021) yields a consistent value of $\pi_{{\rm S},\rm corr}=0.207\pm 0.025\,{\rm mas}$. We fit the spectral energy distribution (SED) using photometric data from 2MASS, Gaia and SkyMapper Southern Survey Data Release 4 (SMSS DR4; Onken et al., 2024). We employ the MESA Isochrones and Stellar Tracks (MIST; Choi et al., 2016) to generate synthetic magnitudes as a function of stellar mass, age, and metallicity. The extinction $E(B-V)$ is taken as a free parameter with a flat prior, and we assume that $R_{V}=3.1$. We use the corrected Gaia parallax as a prior for source distance. MIST returns the stellar radius $R_{*}$, and hence, the angular radius $\theta_{*}=R_{*}/D_{\rm S}$ can be estimated with MCMC. The photometric error bars are inflated by a factor of 2 so that $\chi^{2}/{\rm d.o.f}\sim 1$. We find that the best-fit $E(B-V)=0.64^{+0.06}_{-0.04}$, which is consistent with the lower limit of $E(B-V)>0.61$ (for distance greater than $4.3$ kpc) according to the 3D dust map of Guo et al. (2021). The best-fit stellar parameters are $T_{\rm eff}=3747^{+46}_{-18}\,{\rm K},\,\log g=1.07^{+0.11}_{-0.04}$ and $[{\rm Fe/H}]=0.01^{+0.12}_{-0.07}$, arriving at $\theta_{*}=44.0\pm 0.8\,\mu\rm as$. The SED-derived stellar parameters are consistent with the spectroscopic ones from § 2.3.
[…]
@@ 5.2 Physical Properties of the Lens System
and adopting $\pi_{\rm S}=0.185\pm 0.016\,{\rm mas}$, we can also infer the lens distance,
@@
The lens properties of the favored solution (“wide A--”) are listed in Table 3. The lens system is composed of two M dwarfs with $M_{1}=0.261^{+0.009}_{-0.008}\,M_{\odot}$ and $M_{2}=0.252\pm{0.017}\,M_{\odot}$, separated by a projected separation of $r_{\perp}=7.42\pm 0.33\,{\rm AU}$, and at a distance of $D_{\rm L}=2.31\pm 0.09$ kpc.
Based on MIST isochrones, we estimate that the lens system has $K_{s}=18.5$ (assuming zero extinction). In comparison, the source’s 2MASS magnitude is $K_{s}=9.23$. During the VLTI observations, the magnification is more than $15$, and thus the flux ratio between the lens and magnified source is $\xi=f_{{\rm L}}/(Af_{{\rm S}})<1.4\times 10^{-5}$. Therefore, the lens light has a negligible effect on the closure phase, justifying our approach of calculating closure phases without the lens flux contribution.
@@ 6 Discussion and Conclusion
ASASSN-22av is the first binary-lens microlensing event with multiple images resolved by interferometry. The modeling is considerably more complex than for a single-lens event. For a typical single-lens event, the simple model of two point images is generally satisfactory for fitting the closure-phase data (Dong et al., 2019), and thus, closure-phase modeling can, in principle, be performed separately from the light curve. In contrast, a binary lens has three or five images for a point source, depending on whether the source is outside or inside the caustics. During the VLTI observations of ASASSN-22av, the images, formed from a finite source crossing caustics, morphed from a nearly complete Einstein ring to separated arcs over a period of several nights. Thus, the interpretation of the closure-phase data from such a binary-lens event requires inputs from light-curve modeling.
@@
The weak binary perturbation and sparse pre-peak photometric coverage make the light-curve modeling particularly complicated for ASASSN-22av. In the early stages of our work where we did not include the Gaia light curve, we found 25 local minima in our binary-lens grid search. After incorporating the VLTI data, the best-fit solution, which is the same as the presently-favored “wide A--” solution, ‘predicted’ the existence of the bump at $\rm HJD^{\prime}\sim 9300$ shown in the Gaia light curve. We only realized this ‘prediction’ later, after conducting light-curve modeling after the inclusion of Gaia data. The predictive power of the “wide A--” solution offers good evidence for its validity.
@@
We also made several consistency checks between the light curve and multi-epoch VLTI data. For the favored solution, there are good agreements in multiple aspects, including $\alpha$ (Figure 6), direction of ${\bm{\pi}}_{\rm E}$ (Figure 7), source trajectory (Figure 9) and $\theta_{\rm E}$ (see discussion in § 5.1), further demonstrating the model’s robustness. The agreement between the best-fit model and the closure-phase data is at $\lesssim 0.3^{\circ}$ (Figure 8, §4.3), which likely reflects the instrument’s systematics floor. We measure the Einstein radius $\theta_{\rm E}=0.726$ mas with $0.3\%$ precision, even thought it is close to the smallest value resolvable by VLTI GRAVITY. The expected closure phase would be $\lesssim 5^{\circ}$ for $\theta_{\rm E}$ to $0.6\,{\rm mas}$ and $\lesssim 2^{\circ}$ for $\theta_{\rm E}=0.5\,{\rm mas}$.
@@
The lens system consists of two M dwarfs separated by $\sim 7.4\,{\rm AU}$ at $2.3\,\rm kpc$, with $(M_{1},M_{2})=(0.261\pm{0.009},0.252\pm{0.017})\,M_{\sun}$. The primary mass is more precisely determined than the secondary ($3.4\%$ versus $6.7\%$). This is because, in the wide-binary ($s>1$) regime, the scale of images (measured by VLTI GRAVITY) is proportional to $\theta_{\rm E}$ of the primary mass rather than the total mass. Hence, the uncertainty of the secondary mass $M_{2}$ is contributed by an additional uncertainty from the mass ratio. The estimated orbital period is $P\sim 27\,\rm yr$ for a circular orbit, and the angular separation between the binaries is $\sim 3\,{\rm mas}$ at $2.3\,\rm kpc$. Such faint low-mass binary systems at intermediate separations have so far been inaccessible by any other method (El-Badry, 2024, see their Figure 1). Similar VLTI microlensing observations will be able to identify and characterize intermediate binaries containing dark compact objects like neutron stars and stellar-mass black holes to complement other detection techniques.
@@ Acknowledgment
We thank Cheongho Han and Gregory Green for helpful discussions. This work is based on observations collected at the European Southern Observatory under Director Discretionary Time program 108.23MK. This work is supported by the National Natural Science Foundation of China (Grant No. 12133005) and the science research grants from the China Manned Space Project with No. CMS-CSST-2021-B12. SD acknowledges the New Cornerstone Science Foundation through the XPLORER PRIZE. CSK is supported by NSF grants AST-2307385 and 2407206. This research was funded in part by National Science Centre, Poland, grants OPUS 2021/41/B/ST9/00252 and SONATA 2023/51/D/ST9/00187 awarded to PM.
This research uses data obtained through the Telescope Access Program (TAP), which has been funded by the TAP member institutes. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. We acknowledge the Gaia Photometric Science Alerts Team. We also acknowledge the devoted staff of the Anglo-Australian Telescope who have maintained the HERMES spectrograph over a decade.
This research has made use of the VizieR catalogue access tool, CDS, Strasbourg, France. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation.

### v3

@@ 4.3 Joint Analysis with Light Curves
In this section, we present the joint analysis of VLTI and light-curve modeling. We begin by comparing the light-curve parameters and VLTI constraints. Figure 6 shows the comparisons between the $\alpha$ parameters $\alpha_{\rm LC}$ derived from the light-curve solutions with $\alpha_{\rm VLTI}$. The three solutions are mutually compatible within $3\sigma$, and they are all at $\Psi\sim-100^{\circ}$. These include one close solution (“close A--”) and two wide solutions (“wide A--” and “wide B$+$”). Next we check the consistency of the direction ${\bm{\pi}}_{\rm E}$ between the light-curve solutions and VLTI solution at $\Psi\sim-100^{\circ}$. Figure 7 displays the posterior distribution of $(\pi_{\rm E,E},\pi_{\rm E,N})$ and the $\Phi_{{\bm{\pi}}}$ constraint from VLTI. For three light-curve solutions (“wide A--”, “wide B$+$”, and “close A--”), the two directions agree to $\sim 1\sigma$.
Then we perform a joint fit to the VLTI and light-curve data for the “wide A--”, “wide B$+$”, and “close A--” solutions. We include the higher-order effects of the microlens parallax and binary orbital motion. In addition, we account for the limb-darkening effects in the closure-phase modeling by adopting $\Gamma_{K}=0.29$, based on our adopted source stellar parameters (§ 3.1). Compared to the uniform-brightness profile, limb darkening presents minor improvements in the goodness of fit, with $\Delta\chi^{2}\sim(3,7,8)$ for the solutions (wide A--, wide B$+$, close A--), respectively, and the best-fit parameters with and without the limb-darkening effects are consistent with each other. Thus, the non-limb-darkening approximation adopted in the closure-phase analysis is adequate for probing local minima with grid searches (§ 4.2). Nevertheless, we consider the limb-darkening effects to derive the final results from the joint fit.
@@
Table 2 lists the best-fit parameters and their uncertainties. The global best fit is the “wide A--” solution, and Figure 8 shows its best-fit closure-phase models along with the geometric configurations of the source, caustic, and images. The “close A--” solution is worse by $\Delta\chi^{2}\approx 45$, which is somewhat ($\approx 25$) less compared with the light-curve-only $\Delta\chi^{2}$. Therefore, the preference for the “wide A--” over the “close A--” comes from the light-curve fitting. It mainly stems from the $\sim 0.1$ mag broad bump around $\rm HJD^{\prime}\sim 9300$ in the Gaia light curve, as mentioned in § 3.1. This feature can be well matched by the wide solutions. In contrast, for the “close A--” solutions, the orbital motion parameters can be tuned to induce source crossing of a planetary caustic at $\rm HJD^{\prime}\sim 9300$, but the corresponding perturbation is too sharply peaked to reproduce the observed broad feature (see the right panel of Figure 3). Among the wide solutions, the “wide B$+$” model fit is worse by $\Delta\chi^{2}\approx 296$ with respect to “wide A--,” significantly increased by $\approx 245$ compared with light curve only. As discussed in § 4.2, such $u_{0}>0$ geometries are strongly disfavored by the constraints from multi-epoch VLTI data.
@@
To further understand the VLTI constraints on these three solutions, we performed additional fitting with the interferometric data only. We hold the parameters $(s,q,\rho,\Psi,\theta_{\rm E})$, which define the caustics’ configuration, fixed at their best-fit values from the joint modeling, while the source positions $(u_{x},u_{y})$ of the three nights are allowed to vary. Figure 9 shows the $3\sigma$ $(u_{x},u_{y})$ contours (magenta) compared with the expected source positions (black dots) on the best-fit trajectory (black solid line) from the joint modeling for each solution. For the “wide A--” solution, there is a $\gtrsim 3\sigma$ discrepancy in $(u_{x},u_{y})$ during Night 1, which corresponds to the $\lesssim 0.^{\circ}3$ systematic residuals of the closure phases for U4U2U1 and U4U3U2 in both exposures of that night, as displayed in the top two subpanels of Figure 8. During the other two nights, the source positions are in good agreement with the “wide A--” solution. Note that such $\sim 0.3^{\circ}$ systematic offsets also show up in the third exposure of Night 2 (see Figure 8), but they are not present in the other three exposures (i.e., Night 2-1, 2-2, and 2-4), which are in good agreement with the model. This suggests the existence of sub-degree-level systematic error in the closure-phase data. The “close A--” solution shows a similar degree of consistency to the “wide A--.” In contrast, for the “wide B+” solution, which is ruled out by the VLTI data, the source positions of all three nights deviate by $>3\sigma$ from the expectation.
@@
In conclusion, because the “wide A--” solution is preferred over all other local minima by $\Delta\chi^{2}\gtrsim 45$, we regard it as the only favored solution of ASASSN-22av.
@@ 5.1 Source Properties and Finite-source Effects
The finite-source effects exhibited in ASASSN-22av’s light curves provide an opportunity to cross-check the VLTI-measured $\theta_{\rm E}$ with $\theta_{\rm E}={\theta_{*}}/{\rho}$. The standard method (Yoo et al., 2004) for deriving $\theta_{*}$ for bulge microlensing events uses the source’s dereddened color and magnitude with the extinction coefficients estimated from the location of bulge red clump (RC) stars in a color-magnitude diagram (CMD). In our case, the source star is not toward the bulge, and the CMD contains RC stars at a variety of distances, making it difficult to apply the standard method. Gaia DR3 reports a trigonometric parallax $\pi_{\rm S}=0.137\pm 0.016\,{\rm mas}$. We adopt $\pi_{{\rm S},\rm corr}=0.185\,{\rm mas}$ after applying the preliminary zero-point correction from Lindegren et al. (2021b). The independent investigation on the Gaia parallax zero-point offset by Groenewegen (2021) yields a consistent value of $\pi_{{\rm S},\rm corr}=0.207\pm 0.025\,{\rm mas}$. We fit the spectral energy distribution (SED) using photometric data from 2MASS, Gaia and SkyMapper Southern Survey DR4 (Onken et al., 2024). We employ the MESA Isochrones and Stellar Tracks (MIST; Choi et al., 2016) to generate synthetic magnitudes as a function of stellar mass, age, and metallicity. The extinction $E(B-V)$ is taken as a free parameter with a flat prior, and we assume that $R_{V}=3.1$. We use the corrected Gaia parallax as a prior for source distance. MIST returns the stellar radius $R_{*}$, and hence the angular radius $\theta_{*}=R_{*}/D_{\rm S}$ can be estimated with MCMC. The photometric error bars are inflated by a factor of 2, so that $\chi^{2}/$dof $\sim 1$. We find that the best-fit $E(B-V)=0.64^{+0.06}_{-0.04}$, which is consistent with the lower limit of $E(B-V)>0.61$ (for distance greater than $4.3$ kpc), according to the three-dimensional dust map of Guo et al. (2021). The best-fit stellar parameters are $T_{\rm eff}=3747^{+46}_{-18}\,{\rm K},\,\log g=1.07^{+0.11}_{-0.04}$ and $[{\rm Fe/H}]=0.01^{+0.12}_{-0.07}$, arriving at $\theta_{*}=44.0\pm 0.8\,\mu\rm as$. The SED-derived stellar parameters are consistent with the spectroscopic ones from § 2.3.
For the best-fit “wide A--” solution with $\rho=0.058\pm 0.001$, we find $\theta_{\rm E}=0.759\pm 0.019\,{\rm mas}$, which differs by $1.8\,\sigma$ from the $\theta_{\rm E}=0.724\pm 0.002\,{\rm mas}$ determined by VLTI GRAVITY. In comparison, for the “close A--” solution, where $\rho=0.051^{+0.001}_{-0.001}$, the corresponding $\theta_{\rm E}=0.862\pm 0.023\,{\rm mas}$ deviates from the VLTI GRAVITY value of $\theta_{\rm E}\simeq 0.737\pm 0.002\,{\rm mas}$ by $5.4\sigma$. Note that the $\sim 2\%$ error of our $\theta_{*}$ estimate is likely an underestimate, due to unaccounted-for systematic uncertainties. For example, $\theta_{*}$ inferences from the empirical color/surface brightness relations such as Kervella et al. (2004) generally have errors of $\sim 4\%$$5\%$. In any case, for the “wide A--” solution, $\theta_{\rm E}$ derived from VLTI and finite-source effects are basically consistent.
@@ 5.2 Physical Properties of the Lens System
and adopting $\pi_{\rm S}=0.185\pm 0.016\,{\rm mas}$, we can also infer the lens distance:
@@
The lens properties of the favored solution (“wide A--”) are listed in Table 3. The lens system is composed of two M dwarfs with $M_{1}=0.256^{+0.008}_{-0.006}\,M_{\odot}$ and $M_{2}=0.130\pm{0.007}\,M_{\odot}$, separated by a projected separation of $r_{\perp}=6.83\pm 0.31\,{\rm au}$ and at a distance of $D_{\rm L}=2.29\pm 0.08$ kpc.
Based on MIST isochrones, we estimate that the lens system has $K_{s}=18.5$ (assuming zero extinction). In comparison, the source’s 2MASS magnitude is $K_{s}=9.27$. During the VLTI observations, the magnification is more than $15$, and thus the flux ratio between the lens and magnified source is $\xi=f_{{\rm L}}/(Af_{{\rm S}})<1.4\times 10^{-5}$. Therefore, the lens light has a negligible effect on the closure phase, justifying our approach of calculating closure phases without the lens flux contribution.
@@ 6 Discussion and Conclusion
ASASSN-22av is the first binary-lens microlensing event with multiple images resolved by interferometry. The modeling is considerably more complex than for a single-lens event. For a typical single-lens event, the simple model of two point images is generally satisfactory for fitting the closure-phase data (Dong et al., 2019), and thus closure-phase modeling can, in principle, be performed separately from the light curve. In contrast, a binary lens has three or five images for a point source, depending on whether the source is outside or inside the caustics. During the VLTI observations of ASASSN-22av, the images, formed from a finite source crossing caustics, morphed from a nearly complete Einstein ring to separated arcs over a period of several nights. Thus, the interpretation of the closure-phase data from such a binary-lens event requires inputs from light-curve modeling.
@@
The weak binary-lens perturbation and sparse pre-peak photometric coverage make the light-curve modeling particularly complicated for ASASSN-22av. In the early stages of our work, where we did not include the Gaia light curve, we found 25 local minima in our binary-lens grid search. After incorporating the VLTI data, the best-fit solution, which is the same as the presently favored “wide A--” solution, ‘predicted’ the existence of the bump at $\rm HJD^{\prime}\sim 9300$ shown in the Gaia light curve. We only realized this ‘prediction’ later, after conducting light-curve modeling after the inclusion of Gaia data. The predictive power of the “wide A--” solution offers good evidence for its validity.
@@
We also made several consistency checks between the light curve and multi-epoch VLTI data. For the favored solution, there are good agreements in multiple aspects, including $\alpha$ (Figure 6), direction of ${\bm{\pi}}_{\rm E}$ (Figure 7), source trajectory (Figure 9), and $\theta_{\rm E}$ (see the discussion in § 5.1), further demonstrating the model’s robustness. The agreement between the best-fit model and the closure-phase data is at $\lesssim 0.3^{\circ}$ (Figure 8, §4.3), which likely reflects the instrument’s systematics floor. We measure the Einstein radius $\theta_{\rm E}=0.724$ mas with $0.3\%$ precision, even though it is close to the smallest value resolvable by VLTI GRAVITY. The expected closure phase would be $\lesssim 5^{\circ}$ for $\theta_{\rm E}$ to $0.6\,{\rm mas}$ and $\lesssim 2^{\circ}$ for $\theta_{\rm E}=0.5\,{\rm mas}$.
@@
The lens system consists of two M dwarfs separated by $\sim 6.8\,{\rm au}$ at $2.3\,\rm kpc$, with $(M_{1},M_{2})=(0.256\pm{0.008},0.130\pm{0.007})\,M_{\sun}$. The primary mass is more precisely determined than the secondary ($3.1\%$ versus $5.4\%$). This is because, in the wide-binary ($s>1$) regime, the scale of images (measured by VLTI GRAVITY) is proportional to $\theta_{\rm E}$ of the primary mass rather than the total mass. Hence, the uncertainty of the secondary mass $M_{2}$ is contributed by an additional uncertainty from the mass ratio. The estimated orbital period is $P\sim 28\,\rm yr$ for a circular orbit, and the angular separation between the binaries is $\sim 3\,{\rm mas}$ at $2.3\,\rm kpc$. Such faint low-mass binary systems at intermediate orbital separations have so far been inaccessible by any other method (see, e.g., Figure 1 of El-Badry, 2024). Similar VLTI microlensing observations will be able to identify and characterize intermediate binaries containing dark compact objects like neutron stars and stellar-mass black holes to complement other detection techniques.
@@ Acknowledgment
We thank Cheongho Han and Gregory Green for helpful discussions. This work is based on observations collected at the European Southern Observatory under Director Discretionary Time program 108.23MK. This work is supported by the National Natural Science Foundation of China (grant No. 12133005) and the science research grants from the China Manned Space Project with No. CMS-CSST-2021-B12. S.D. acknowledges the New Cornerstone Science Foundation through the XPLORER PRIZE. C.S.K. is supported by NSF grants AST-2307385 and 2407206. This research was funded in part by the National Science Centre, Poland, and grants OPUS 2021/41/B/ST9/00252 and SONATA 2023/51/D/ST9/00187 awarded to P.M.
This research uses data obtained through the Telescope Access Program (TAP), which has been funded by the TAP member institutes. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC; https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. We acknowledge the Gaia Photometric Science Alerts Team. We also acknowledge the devoted staff of the Anglo-Australian Telescope who have maintained the HERMES spectrograph over a decade.
This research has made use of the VizieR catalog access tool, CDS, Strasbourg, France. This publication makes use of data products from the Two Micron All Sky Survey, which is a joint project of the University of Massachusetts and the Infrared Processing and Analysis Center/California Institute of Technology, funded by the National Aeronautics and Space Administration and the National Science Foundation.